OpenStudy (anonymous):

what are the factors of x^2+5x-5

OpenStudy (anonymous):

Hint: You need to use imaginary numbers

OpenStudy (adamaero):

(x )( ) Did you start like that?

OpenStudy (adamaero):

because it wouldn't work; try the quadratic equation

OpenStudy (anonymous):

ok thanks ik how to get it but im not the best at factoring. just needed confirmation

OpenStudy (trojanpoem):

a = 1 b= 5 c = -5 \[\frac{ -b \pm \sqrt{b ^{2}- 4 ac} }{ 2a }\]

OpenStudy (trojanpoem):

\[\frac{ -5\pm \sqrt{25 - 4 * 1 * -5} }{ 2 * 1 }\] \[\frac{ -5\pm \sqrt{45} }{ 2}\]

OpenStudy (trojanpoem):

45 = 5 * 9 = \[\sqrt{9} * \sqrt{5}\] = \[3\sqrt{5}\] \[the factorys are : \frac{ -5 \pm 3\sqrt{5} }{ 2 }\]